At what pressure, in atmospheres, would the number of molecules per unit volume in an ideal gas be equal to the number density of the conduction electrons in copper, with both gas and copper at temperature T = 400 K?
Answer #1

density of the conduction electrons in copper = 1.1*10^29 m^-3

for ideal gas, Apply P*V = n*R*T

P/(R*T) = n/V


P/(8.314*400) = 1.1*10^29

P = 1.1*10^29*8.314*400

= 3.658*10^32 pa

= 3.658*10^32/(1.013*10^5) atm

= 3.61*10^27 atm <<<<<<<<<<<----------Answer